OpenStudy (onepieceftw):

How many liters of oxygen gas, at standard temperature and pressure, will react with 35.4 grams of calcium metal? No idea how to solve this. :/

OpenStudy (anonymous):

2 Ca . . . + . . . O2 . . . -----> . . . 2 CaO . . . . . 40,08 g/mol n(Ca) = m(Ca) / M(Ca) n(Ca) = 35,4 / 40,08 n(Ca) = 0,8832 mol of Ca 2 mol of Ca requires 1 mol of O2 . . . then : 0,8832 mol of Ca requires 0,8832 / 2 = 0,4416 mol of O2 V(O2) = Vm x n(O2) V(O2) = 22,4 x 0,4416 V(O2) = 9,89 L of O2